# Adversarial Training for Failure-Sensitive User Simulation in Mental Health Dialogue Optimization

### v1

@@ 3.3 Base Simulator Training
We begin by training a base user simulator through supervised fine-tuning (SFT) on anonymized user conversations. Starting from Llama-3.3-70B-Instruct as our pre-trained foundation model, we fine-tune on a dataset $\mathcal{D}=\{(h_{i},c_{i},u_{i})\}_{i=1}^{N}$ of real user responses paired with conversation history and context:
@@ 4.1 Domain and Data
Our dataset consists of anonymized real user conversations with our mental health support chatbot, including session-level context (user history, previous session summaries) and turn-level annotations of user actions (e.g., expressing distress, gaining insight, agreeing with suggestions) and system issues (e.g., invalidating responses, hallucinations, unclear wording).
@@ 4.2 Model Configurations
We train simulators by fine-tuning Llama-3.3-70B-Instruct using LoRA (hu2022lora) on datasets of varying sizes (10K, 20K, 40K, 80K, 160K sessions) to explore data scale effects. We perform adversarial training on the final checkpoint (160K), producing UserSim-160K-it1, UserSim-160K-it2, and UserSim-160K-it3 through successive refinement iterations. For comparison, we evaluate three zero-shot base models: GPT-4o, Kimi-K2-Instruct, and Llama-3.3-70B-Instruct. Training hyperparameters are provided in Appendix C.
@@ 7 Conclusion
All data used in this research were collected from real user conversations with our mental health chatbot with explicit user consent. All personally identifiable information was removed through anonymization procedures in accordance with data protection policies and ethical guidelines for human subjects research. However, due to the sensitive nature of mental health data, we are unable to share the datasets used in this research, despite our desire to promote reproducibility and open science.
Our simulators are designed for system evaluation and development, enabling rapid iteration without exposing users to suboptimal experiences. We note that simulators trained on real conversations may perpetuate biases present in the original dataset; applications should include appropriate oversight and ethical review. Additionally, while simulators can reduce the need for human testing in many scenarios, they should complement rather than entirely replace human feedback for validating novel interventions or edge cases.
@@ Appendix A Context Construction for Mental Health Support Chatbot
Current Session Summary: A summary of the current session being simulated, extracted from the original conversation. This grounds the simulator in the specific topics and emotional tone of the target session, reducing aimless or disengaged responses. This session-specific grounding is critical for the simulator to cluster tightly with corresponding real sessions.
@@ B.2 Issue Category Distribution
Figure 4 shows the discrepancy between simulated and real session issue category distributions. The simulator over-represents certain categories, including “Chatbot contradicts itself” ($\sim$1.5 percentage points), “Chatbot forgets in context” ($\sim$1.25 percentage points), and “unclear or confusing wording”—likely because simulated users engage more persistently with problematic chatbot behaviors, triggering repeated issues. Conversely, the simulator under-represents categories like “duplicate chatbot messages” ($\sim$-2.5 percentage points), “pushing techniques after user request to stop” ($\sim$-2.5 percentage points), and “invalidating or dismissive framing,” which require specific conversational histories that may be less common in simulation. The relatively small magnitude of most discrepancies (within $\pm$2.5 percentage points) indicates reasonable distributional alignment overall.
@@ B.5 A/B Test Replication
Appendix C Training Hyperparameters

### v2

@@ 3.3 Base Simulator Training
We begin by training a base user simulator through supervised fine-tuning (SFT) on anonymized user conversations. Starting from Llama-3.3-70B-Instruct (grattafiori2024llama3) as our pre-trained foundation model, we fine-tune on a dataset $\mathcal{D}=\{(h_{i},c_{i},u_{i})\}_{i=1}^{N}$ of real user responses paired with conversation history and context:
@@ 4.1 Domain and Data
Our dataset consists of anonymized real user conversations in English with our mental health support chatbot, including session-level context (user history, previous session summaries) and turn-level annotations of user actions (e.g., expressing distress, gaining insight, agreeing with suggestions) and system issues (e.g., invalidating responses, hallucinations, unclear wording).
@@ 4.2 Model Configurations
We train simulators by fine-tuning Llama-3.3-70B-Instruct using LoRA (hu2022lora) on datasets of varying sizes (10K, 20K, 40K, 80K, 160K sessions) to explore data scale effects. We perform adversarial training on the final checkpoint (160K), producing UserSim-160K-it1, UserSim-160K-it2, and UserSim-160K-it3 through successive refinement iterations. For comparison, we evaluate three zero-shot base models: GPT-4o (openai2024gpt4o), Kimi-K2-Instruct (kimiteam2025k2), and Llama-3.3-70B-Instruct. Training hyperparameters are provided in Appendix C.2.
@@ 5 Results
All reported results are from single training runs per configuration.
@@ 7 Conclusion
Assumptions and Robustness: Our methodology relies on several key assumptions: (1) discriminator-based rewards accurately reflect true realism rather than exploitable artifacts, (2) the 14 chatbot models used for evaluation are representative of the broader space of possible configurations, and (3) GPT-5-based issue detection accurately identifies problematic conversational features with sufficient reliability to validate simulator quality. Results may be sensitive to violations of these assumptions.
@@
All data used in this research were collected from real user conversations with our mental health chatbot with explicit user consent. Personally identifiable information was removed through automated script-based detection and removal of names, phone numbers, email addresses, and other identifiers in accordance with data protection policies and ethical guidelines for human subjects research. However, due to the sensitive nature of mental health data, we are unable to share the datasets used in this research, despite our desire to promote reproducibility and open science.
Realistic user simulators could potentially be misused to train systems optimized for engagement over user wellbeing, or to manipulate vulnerable populations. Our simulators are designed for system evaluation and development, enabling rapid iteration without exposing users to suboptimal experiences. We note that simulators trained on real conversations may perpetuate biases present in the original dataset, potentially disadvantaging certain demographic groups. We emphasize that simulators should be deployed with appropriate oversight and ethical review. Additionally, while simulators can reduce the need for human testing in many scenarios, they should complement rather than entirely replace human feedback for validating novel interventions or edge cases.
Study procedures were approved as exempt by the Institutional Review Board at New York University School of Medicine (i25-01177).
@@ Appendix A Context Construction for Mental Health Support Chatbot
Current Session Summary: A summary of the current session being simulated, extracted from the original conversation. This grounds the simulator in the specific topics and emotional tone of the target session, reducing aimless or disengaged responses.
@@ B.2 Issue Category Distribution
Figure 4 shows the discrepancy between simulated and real session issue category distributions. The simulator over-represents certain categories, including “Chatbot contradicts itself”, “Chatbot forgets in context”, and “Unclear or confusing wording”—likely because simulated users engage more persistently with problematic chatbot behaviors, triggering repeated issues. Conversely, the simulator under-represents categories like “Duplicate chatbot messages”, “Chatbot hallucinates”, and “Premature wrap-up”, which require specific conversational histories that may be less common in simulation. The relatively small magnitude of most discrepancies (within $\pm$2.5 percentage points) indicates reasonable distributional alignment overall.
@@ B.5 A/B Test Replication
Appendix C Model and Training Details
C.1 Model Licenses and Terms of Use
All models used in this work are distributed under licenses permitting research use. Llama-3.3-70B-Instruct and Llama-3.1-8B-Instruct are released under the Llama 3.3 and Llama 3.1 Community License Agreements respectively, which permit commercial and research use with certain restrictions (e.g., services with over 700 million monthly active users require additional licensing). GPT-4o is accessed via OpenAI’s API under their Service Terms, which grant users ownership of model outputs for both research and commercial applications. Kimi-K2-Instruct is released under a modified MIT license that requires prominent display of the model name when generating content commercially.
C.2 Training Hyperparameters